OpenStudy (anonymous):

how do you rationalize the denominator for 4/ (1+ sqrt 3)

OpenStudy (anonymous):

multiply it with (1-sqrt3)/(1-sqrt3)

OpenStudy (anonymous):

so \[4+\sqrt 3 \div 1+\sqrt 3\] ?

OpenStudy (anonymous):

the denominator you wrote must be -2

OpenStudy (anonymous):

so.... \[4+\sqrt 3 \div -2\]

OpenStudy (anonymous):

so what is wrong with that? you just did it!

OpenStudy (anonymous):

awesome! thanks!

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

to simplify that would you divide 4/-2? or is it already in simplest form

OpenStudy (anonymous):

yes u can divide it it would become -2-(sqrt 3)/2=-2.865

OpenStudy (anonymous):

ok :)